# Natural Fingerprints of Large Language Models

### v1

@@ Abstract
Large language models (LLMs) often exhibit biases—systematic deviations from expected norms—in their outputs. These range from overt issues, such as unfair responses, to subtler patterns that can reveal which model produced them. We investigate the factors that give rise to identifiable characteristics in LLMs. Since LLMs model training data distribution, it is reasonable that differences in training data naturally lead to the characteristics. However, our findings reveal that even when LLMs are trained on the exact same data, it is still possible to distinguish the source model based on its generated text. We refer to these unintended, distinctive characteristics as natural fingerprints. By systematically controlling training conditions, we show that the natural fingerprints can emerge from subtle differences in the training process, such as parameter sizes, optimization settings, and even random seeds. We believe that understanding natural fingerprints offers new insights into the origins of unintended bias and ways for improving control over LLM behavior.
@@ 1 Introduction
Recent studies have shown that large language models (LLMs) display distinctive characteristics in their outputs, enabling accurate identification of which specific model generated a given text (Antoun et al., 2023; Sun et al., 2025). However, the underlying causes of these distinctive patterns are under-explored. The characteristics of LLMs are primarily determined by their training data design, but unintended biases can also emerge. LLMs may demonstrate skewed perspectives toward certain demographic groups (Gallegos et al., 2024) or produce unfair assessments when employed as evaluative tools (LLM-as-a-Judge) (Zheng et al., 2023; Chen et al., 2024) due to the unintended biases. Understanding what determines the characteristics of LLM outputs is thus crucial for addressing these issues and controlling LLM behavior.
@@
In this work, we explore the sources of distinctive patterns in LLM outputs using an approach inspired by dataset classification (Torralba & Efros, 2011; Liu & He, 2024). In the original setting, the task involves identifying which dataset a given data instance comes from, and high classification accuracy reveals hidden, often unintended, signatures specific to each dataset. Analogously, we classify text samples by their source LLM, as shown in Figure 1. Reliable identification of model’s outputs implies the existence of model-specific patterns. This allows us to probe the underlying causes of these patterns and better understand LLM behavior. Specifically, we begin by testing whether outputs from publicly available pretrained models can be distinguished through classification. We then train models from scratch under controlled settings, gradually eliminating differences in parameter size and training configurations, and examine whether the source model can still be identified from its output. Thorough the experiments, we isolate the factors that contribute to these model-specific patterns.
@@
Our experiments revealed several findings: the outputs of publicly available LLMs can be classified with high accuracy. Since they are trained on different datasets, this result indicates the presence of dataset-specific patterns in their outputs, which is reasonable given that LLMs learn to model the distribution of their training data. In addition, surprisingly, even LLMs trained on the same dataset can be distinguished above chance rate. By systematically controlling training conditions, we found that hyperparameters such as learning rate, weight decay, and even random seeds contribute to the emergence of model-specific signatures. In particular, the training data order has a relatively strong influence on the characteristics of LLM outputs.
Unlike deliberate model fingerprinting methods (Uchida et al., 2017; Gu et al., 2022; Xu et al., 2024; Yamabe et al., 2024), the patterns we observe emerge spontaneously during the training process without explicit design—a characteristic we term natural fingerprints. The natural fingerprints can arise from extremely subtle differences, such as the order in which training data is presented, indicating that it is difficult to completely eliminate them in practical model development. Therefore, we believe that understanding them is essential for interpreting differences between models beyond performance alone.
@@ 2.1 Bias in Large Language Models
LLMs consistently exhibit various forms of “bias”, referring to systematic deviations from expected or intended norms in their outputs. These biases manifest in multiple ways: normative biases, which result in unfair treatment of certain demographic groups (Gallegos et al., 2024), and evaluative biases, where LLMs used as judges display systematic preferences toward particular models or content types (Zheng et al., 2023; Chen et al., 2024). Such biases are natural consequences of training data distributions or architecture’s inductive biases, often unintended by developers.
@@ 2.2 LLM Output Identification
Identifying LLM-generated content has become a topic of practical interest. Distinguishing machine-generated text from human-written text is crucial for detecting the misuse of text generation models (Gehrmann et al., 2019; Zellers et al., 2019b; Fagni et al., 2021; Solaiman et al., 2019; Guo et al., 2023; Mitrović et al., 2023; Mitchell et al., 2023; Sadasivan et al., 2023). Some research focuses on watermarking, deliberately embedding detectable traces into model outputs (Kirchenbauer et al., 2023; Kuditipudi et al., 2023; Christ et al., 2023; Yang et al., 2023), or fingerprinting, identifying the model itself (Uchida et al., 2017; Gu et al., 2022; Xu et al., 2024; Yamabe et al., 2024). In contrast, our work addresses the fundamental question of how model-specific traces naturally emerge during training.
Among the closest related works, Antoun et al. (2023) first demonstrated that outputs from different LLMs can be distinguished through classification. Sun et al. (2025) analyzed several commercial LLMs to identify distinguishing features in their outputs, finding that word choice, formatting conventions, and semantic patterns serve as recognizable characteristics. Our research extends these findings by first confirming the distinguishability of public LLMs and then training controlled models to systematically investigate the factors contributing to these natural fingerprints.
@@ 3.1 From Dataset Classification to LLM Classification
Dataset classification is a task that involves identifying which dataset an image belongs to. Originally proposed by Torralba & Efros (2011), this experiment highlights unintended biases in datasets. If classification is highly accurate, it suggests that datasets contain unique characteristics beyond their intended content, such as differences in quality, acquisition methods, or styles.
We redesign the dataset classification experiment as an LLM classification task to analyze biases in LLMs. Formally, let =$\mathcal{M}=\{M_{1},M_{2},\ldots,M_{m}\}$ be a set of $m$ distinct language models, and $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{n}\}$ be a set of $n$ prompts. For each prompt $p_{i}\in\mathcal{P}$ and model $M_{j}\in\mathcal{M}$, we obtain a generated text $t_{i,j}=M_{j}(p_{i})$ where $M_{j}(p_{i})$ denotes the output of model $M_{j}$ given prompt $p_{i}$. We then construct a dataset $\mathcal{D}=\{(t_{i,j},j)\mid i\in\{1,2,\ldots,n\},j\in\{1,2,\ldots,m\}\}$ of pairs, where each pair consists of a generated text and the index of its source model.
@@ 3.2 Sampling Texts from LLMs
As prompts $\mathcal{P}$, we use the first 50 characters from one million web pages sampled from the CommonCrawl subset of SlimPajama (Soboleva et al., 2023). Each language model $M_{j}$ generates continuations from the prompts using pure sampling, without adjusting decoding parameters. The maximum number of output tokens was set to 512. To filter out degenerate text, we removed samples where any 5-gram appeared 8 or more times. Note that in the resulting classification dataset, the text $t_{i,j}$ is the concatenation of the prompt and the LLM’s continuation.
To prevent data leakage between splits, we remove near-duplicates using semantic deduplication (Abbas et al., 2023) with an English embedding model [^1] https://huggingface.co/thenlper/gte-small. With the similarity threshold set at 0.2, this process removed about 0.1% of the data.
@@ 3.3 Classification Setup
3.3.1 Unigram-based Classifier
@@
We use LIBLINEAR (Fan et al., 2008) for implementation. We tokenize datasets with the Llama2 tokenizer (Touvron et al., 2023a) and use unigram frequencies to construct feature vectors. We scale each feature value from 0 to 1.
3.3.2 Transformer-based Classifier
We also perform classification with a pretrained Transformer, DeBERTaV3 (He et al., 2021), with a linear classification head. In this setup, we verify whether the text contains higher-level patterns beyond word frequencies that indicate LLM-specific biases.
For training, we use the Adam optimizer (Kingma & Ba, 2014) with a batch size of 64 and a peak learning rate of 2e-5 for 1 epochs. We employ early stopping based on accuracy on the validation set. The learning rate is linearly warmed up in the first 10% of the training steps and then linearly decayed to 0 over the remaining training period.
@@ 4.1 Distinguishing Between Different LLMs
First, we use LLMs with approximately the same number of parameters but trained on different datasets and under different conditions. We choose six pretrained LLM families publicly available at Hugging Face Hub: Llama2 (Touvron et al., 2023b), Mistral (Jiang et al., 2023), Qwen2 (Yang et al., 2024), Falcon (Almazrouei et al., 2023), Pythia (Biderman et al., 2023), and Sarashina2 (SB Intuitions, 2024). We chose these models because their training data reportedly contains substantial amounts of English text, while still maintaining diversity (Qwen2 incorporates Chinese corpus, while Sarashina2 includes Japanese).
@@
Model Unigram feature
Llama2 Label _… _, _. _ The An In _? Tags
[…]
Qwen2 • By ^M NUMBER In <0xE6> _\$ , _. The 的
Falcon << .. ,. _Share Abstract IMAGE -. - Contents ”
[…]
Sarashina2 yy … © XXXX xxxx xxx _- erved NAME Top
[…]
Table 1: The top 10 largest weighted features to detect each LLM. We represent a whitespace with _—.
@@
While these features indicate the influence of training data on LLM bias, the models also differ in other aspects such as architecture and training hyperparameters. Which specific differences give rise to the distinctive patterns remains under-explored, although the fact that outputs from LLMs trained on different datasets can be classified has already been established in the previous studies Antoun et al. (2023); Sun et al. (2025). In the following experiments, we systematically eliminate each factor by training models from scratch under controlled conditions to explore causes of bias.
@@ 4.2 Removing the Influence of Training Data
First, we train three models with different parameter sizes, 0.5B, 1B, and 3B, with the same training corpus. The training corpus consists of 100B tokens extracted from FineWeb-Edu (Penedo et al., 2024), tokenized with the Llama2 tokenizer. We use Megatron-LM (Shoeybi et al., 2020) for training. Other hyperparameters are described in Appendix A.
@@
However, the classification accuracy still exceeds chance rate, indicating that factors beyond training data differences also contribute to bias. Model size may be a major factor here, but this effect can be decomposed into two aspects: how well the model absorbs the training data distribution, and differences arising from weight initialization across varying parameter spaces. As our later experiments show, even when controlling for parameter size, initialization alone can lead to distinguishable outputs (Section 4.4), suggesting different ways model-specific patterns can develop.
@@ 4.3 Removing the Influence of Parameter Sizes
Next, we evaluate whether it is possible to classify generated text when models with identical architecture (including number of layers and hidden sizes) are trained on the same dataset. This allows us to assess the extent to which factors other than architecture contribute to model-specific biases.
Specifically, we train seven 1B parameter LLMs with different learning rates, weight decay values, and random seeds. We confirmed that the trained models have comparable capabilities by verifying their downstream scores are nearly identical (Appendix C).
@@
Compared to the previous experiment, the Transformer-based classification accuracy decreased, suggesting that differences in model size (e.g., particularly in the capacity to capture the training data distribution) may contribute to the emergence of higher-order distinguishable patterns beyond simple word frequency differences.
@@ 4.4 What kind of Randomness during Training Induces Bias in LLMs?
To further investigate how these “natural fingerprints” develop throughout the training process, we examined whether model identifiability evolves predictably over time. We conducted additional experiments with models trained on 1T tokens instead of 100B tokens. We trained three different Transformer-based classification models under both the Order setting (different training data order, same initialization) and the Init setting (different initialization, same training data orders), and performed classification experiments at six training checkpoints: 10B, 50B, 100B, 200B, 500B, 1T tokens (Figure 3).
@@
This convergent behavior suggests that despite differences in optimization paths—whether due to initialization or data order—models tend to exhibit their behavior when trained sufficiently on the same underlying dataset. However, in practical model development, it is often infeasible to perfectly control all training conditions except for the random seed and train models for sufficient steps. In other words, completely eliminating natural fingerprints is unlikely in real-world settings. Fortunately, the relatively small gap between accuracy of unigram-based and Transformer-based classifiers in Table 4 suggests that, when training data and model size are controlled, the distinguishability may largely be attributed to differences in word frequency distributions, which are unlikely to cause significant practical issues.
4.5 SFT experiment
Finally, we test whether natural fingerprints also appear in smaller-scale training, such as supervised fine-tuning (SFT). In this setting, the model starts from pretrained weights, which can impose stronger effects to the training dynamics compared to the random weights in the pretraining experiments.
We fine-tune the 1B pretrained models (trained on 100B tokens in the previous experiment) using the Alpaca dataset (Taori et al., 2023). We use a learning rate of 2e-5 and train for one epoch with AdamW. Other hyperparameters are listed in Appendix B.

### v2

@@ Abstract
Recent studies have shown that the outputs from large language models (LLMs) can often reveal the identity of their source model. While this is a natural consequence of LLMs modeling the distribution of their training data, such identifiable traces may also reflect unintended characteristics with potential implications for fairness and misuse. In this work, we go one step further and show that even when LLMs are trained on exactly the same dataset, their outputs remain distinguishable, suggesting that training dynamics alone can leave recognizable patterns. We refer to these unintended, distinctive characteristics as natural fingerprints. By systematically controlling training conditions, we show that the natural fingerprints can emerge from subtle differences in the training process, such as parameter sizes, optimization settings, and even random seeds. These results suggest that training dynamics can systematically shape model behavior, independent of data or architecture, and should be explicitly considered in future research on transparency, reliability, and interpretability.
@@ 1 Introduction
Recent studies have shown that large language models (LLMs) display distinctive characteristics in their outputs, enabling accurate identification of the source models (Antoun, Sagot, and Seddah 2023; Sun, Yin et al. 2025; Gao, Liang, and Guestrin 2024). These results are largely based on experiments using LLMs trained on different datasets, where such differences naturally lead to identifiable patterns due to the models’ role in modeling data distributions. However, it remains unclear whether the distinguishability of model-generated texts arises solely from dataset variations or if other subtle factors within the training procedure itself also contribute to such characteristics. LLMs may demonstrate skewed perspectives toward certain demographic groups (Gallegos, Rossi et al. 2024) or produce unfair assessments when employed as evaluative tools (LLM-as-a-Judge) (Zheng, Chiang et al. 2023; Chen, Chen et al. 2024) due to the unintended biases. Deeply understanding what determines the characteristics of LLM outputs is thus crucial for addressing these issues and controlling LLM behavior.
@@
In this paper, we extend this line of research by examining subtle variations in the training process, even in scenarios where models are trained using exactly the same datasets, to determine whether these differences create identifiable model-specific characteristics. For this purpose, we adopt dataset classification (Torralba and Efros 2011; Liu and He 2024), which involves identifying which dataset a given data instance comes from, and high classification accuracy reveals hidden, often unintended, signatures specific to each dataset. Analogously, we classify text samples by their source LLM, as shown in Figure 1. Reliable identification of model’s outputs implies the existence of model-specific patterns. This allows us to probe the underlying causes of these patterns and better understand LLM behavior. Specifically, we begin by testing whether outputs from publicly available pretrained models can be distinguished through classification as a baseline. We then train models from scratch under controlled settings, gradually eliminating differences in parameter size and training configurations, and examine whether the source model can still be identified from its output. Thorough the experiments, we isolate the factors that contribute to these model-specific patterns.
@@
Our experiments revealed several findings: even when LLMs are trained on the same dataset, their outputs can still be distinguished above chance rate. While it is already known that LLMs trained on different datasets produce identifiable outputs (Antoun, Sagot, and Seddah 2023; Gao, Liang, and Guestrin 2024; Sun, Yin et al. 2025), our results show that training dynamics alone can give rise to model-specific signatures. By systematically controlling training conditions, we found that hyperparameters such as learning rate, weight decay, and even random seeds contribute to the emergence of model-specific signatures. In particular, the training data order has a relatively strong influence on the characteristics of LLM outputs.
Unlike deliberate model fingerprinting methods (Uchida, Nagai et al. 2017; Gu, Huang et al. 2022; Xu, Wang et al. 2024; Yamabe, Takahashi et al. 2024), the patterns we observe emerge spontaneously during the training process without explicit design—a characteristic we term natural fingerprints. In doing so, our study introduces a new axis of identifiability in LLMs. This suggests that future research on transparency, reliability, and interpretability should not only consider the influence of training data, but also account for the procedural aspects of model training.
@@ 2.1 Bias in Large Language Models
LLMs consistently exhibit various forms of “bias”, referring to systematic deviations from expected or intended norms in their outputs. These biases manifest in multiple ways: normative biases, which result in unfair treatment of certain demographic groups (Gallegos, Rossi et al. 2024), and evaluative biases, where LLMs used as judges display systematic preferences toward particular models or content types (Zheng, Chiang et al. 2023; Chen, Chen et al. 2024). Such biases are natural consequences of training data distributions or architecture’s inductive biases, often unintended by developers.
@@ 2.2 LLM Output Identification
Identifying LLM-generated content has become a topic of practical interest. Distinguishing machine-generated text from human-written text is crucial for detecting the misuse of text generation models (Gehrmann, Strobelt, and Rush 2019; Zellers et al. 2019; Fagni, Falchi et al. 2021; Solaiman, Brundage et al. 2019; Guo, Zhang et al. 2023; Mitrović, Andreoletti, and Ayoub 2023; Mitchell, Lee et al. 2023; Sadasivan, Kumar et al. 2023). Some research focuses on watermarking, deliberately embedding detectable traces into model outputs (Kirchenbauer, Geiping et al. 2023; Kuditipudi, Thickstun et al. 2023; Christ, Gunn, and Zamir 2023; Yang, Chen et al. 2023), or fingerprinting, identifying the model itself (Uchida, Nagai et al. 2017; Gu, Huang et al. 2022; Xu, Wang et al. 2024). In contrast, our work addresses the fundamental question of whether model-specific traces naturally emerge during training.
Among the closest related works, Antoun, Sagot, and Seddah (2023) first demonstrated that outputs from different LLMs can be distinguished through classification. Gao, Liang, and Guestrin (2024) proposed Model Equality Testing, a statistical framework for distinguishing black-box LLMs based on their output distributions. Sun, Yin et al. (2025) analyzed several commercial LLMs to identify distinguishing features in their outputs, finding that word choice, formatting conventions, and semantic patterns serve as recognizable characteristics.
While previous studies have examined the identifiability of LLM outputs using models trained on different datasets (Antoun, Sagot, and Seddah 2023; Gao, Liang, and Guestrin 2024; Sun, Yin et al. 2025), their experimental setups inevitably conflate the effects of data variation with other factors. In contrast, our work isolates the impact of training configurations by holding the dataset constant and varying elements such as initialization, data order, and hyperparameters. Our findings reveal that these seemingly minor differences can result in consistent and detectable patterns in model outputs.
@@ 3.1 From Dataset Classification to LLM Classification
Dataset classification is a task that involves identifying which dataset an image belongs to. Originally proposed by Torralba and Efros (2011), this experiment highlights unintended biases in datasets. If classification is highly accurate, it suggests that datasets contain unique characteristics beyond their intended content, such as differences in quality, acquisition methods, or styles.
We redesign the dataset classification experiment as an LLM classification task to analyze biases in LLMs. Formally, let $\mathcal{M}=\{M_{1},M_{2},\ldots,M_{m}\}$ be a set of $m$ distinct language models, and $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{n}\}$ be a set of $n$ prompts. For each prompt $p_{i}\in\mathcal{P}$ and model $M_{j}\in\mathcal{M}$, we obtain a generated text $t_{i,j}=M_{j}(p_{i})$ where $M_{j}(p_{i})$ denotes the output of model $M_{j}$ given prompt $p_{i}$. We then construct a dataset $\mathcal{D}=\{(t_{i,j},j)\mid i\in\{1,2,\ldots,n\},j\in\{1,2,\ldots,m\}\}$ of pairs, where each pair consists of a generated text and the index of its source model.
@@ 3.2 Sampling Texts from LLMs
As prompts $\mathcal{P}$, we use the first 50 characters from one million web pages sampled from the CommonCrawl subset of SlimPajama (Soboleva, Al-Khateeb et al. 2023). Each language model $M_{j}$ generates continuations from the prompts using pure sampling, without adjusting decoding parameters. The maximum number of output tokens was set to 512. To filter out degenerate text, we removed samples where any 5-gram appeared 8 or more times. Note that in the resulting classification dataset, the text $t_{i,j}$ is the concatenation of the prompt and the LLM’s continuation.
To prevent data leakage between splits, we remove near-duplicates using semantic deduplication (Abbas, Tirumala et al. 2023) with an English embedding model [^1] https://huggingface.co/thenlper/gte-small. With the similarity threshold set at 0.2, this process removed about 0.1% of the data.
@@ 3.3 Classification Setup
Unigram-based Classifier
@@
We use LIBLINEAR (Fan, Chang et al. 2008) for implementation. We tokenize datasets with the Llama2 tokenizer (Touvron, Martin et al. 2023) and use unigram frequencies to construct feature vectors. We scale each feature value from 0 to 1.
Transformer-based Classifier
We also perform classification with a pretrained Transformer, DeBERTaV3 (He, Gao, and Chen 2021), with a linear classification head. In this setup, we verify whether the text contains higher-level patterns beyond word frequencies that indicate LLM-specific biases.
For training, we use the Adam optimizer (Kingma and Ba 2014) with a batch size of 64 and a peak learning rate of 2e-5 for 1 epochs. We employ early stopping based on accuracy on the validation set. The learning rate is linearly warmed up in the first 10% of the training steps and then linearly decayed to 0 over the remaining training period.
@@ 4.1 Distinguishing Between Different LLMs
First, we use LLMs with approximately the same number of parameters but trained on different datasets and under different conditions. Although prior work has shown that LLMs trained on different datasets produce distinguishable outputs (Antoun, Sagot, and Seddah 2023; Gao, Liang, and Guestrin 2024; Sun, Yin et al. 2025), we begin by replicating this phenomenon in our setting. This serves two purposes: it validates our experimental setup, and it provides a baseline against which we can compare cases where training data is held constant.
We choose six pretrained LLM families publicly available at Hugging Face Hub: Llama2 (Touvron, Martin et al. 2023), Mistral (Jiang, Sablayrolles et al. 2023), Qwen2 (Yang, Yang et al. 2024), Falcon (Almazrouei, Alobeidli et al. 2023), Pythia (Biderman, Schoelkopf et al. 2023), and Sarashina2 (SB Intuitions 2024). We chose these models because their training data reportedly contains substantial amounts of English text, while still maintaining diversity (Qwen2 incorporates Chinese corpus, while Sarashina2 includes Japanese).
@@
Table 1: The top 10 largest weighted features to detect each LLM. We represent a whitespace with _.
@@
While these features indicate the influence of training data on LLM bias, the models also differ in other aspects such as architecture and training hyperparameters. Which specific differences give rise to the distinctive patterns remains under-explored, although the fact that outputs from LLMs trained on different datasets can be classified has already been established in the previous studies (Antoun, Sagot, and Seddah 2023; Sun, Yin et al. 2025). In the following experiments, we systematically eliminate each factor by training models from scratch under controlled conditions to explore causes of bias.
@@ 4.2 Removing the Influence of Training Data
First, we train three models with different parameter sizes, 0.5B, 1B, and 3B, with the same training corpus from scratch. The training corpus consists of 100B tokens extracted from FineWeb-Edu (Penedo, Kydlíček et al. 2024), tokenized with the Llama2 tokenizer. We use Megatron-LM (Shoeybi, Patwary et al. 2020) for training and it takes 800 GPU hours with a H100 GPU for training a 1B model. Other hyperparameters are described in Appendix A.
@@
However, the classification accuracy still exceeds chance rate, indicating that factors beyond training data differences also contribute to distinguishability. Model size may be a major factor here, but this effect can be decomposed into two aspects: how well the model absorbs the training data distribution, and differences in the behavior of models at initialization across varying parameter spaces. As our later experiments (Section 4.4) show, even when controlling for parameter size, initialization alone can lead to distinguishable outputs, but the accuracy decreases compared to Table 2. This suggests that the modeling capability induced by the difference in the parameter size somewhat contributes to distinguishability.
Hyperparameter
lr wd seed PIQA HellaSwag WinoGrande ARC-e ARC-c OBQA
5e-4 0.1 1 71.6 52.2 54.1 63.4 35.0 37.6
5e-4 0.1 2 70.8 51.4 54.1 61.7 34.6 36.0
5e-4 0.1 3 71.0 52.0 54.3 62.2 34.0 39.6
1e-3 0.1 1 71.5 53.8 55.5 64.8 36.4 38.4
1e-4 0.1 1 69.9 46.4 53.9 59.7 31.8 35.6
5e-4 0.01 1 70.4 50.5 52.2 60.7 33.9 37.4
5e-4 0 1 70.7 51.2 54.4 62.2 34.2 37.8
@@ 4.3 Removing the Influence of Parameter Sizes
Next, we evaluate whether it is possible to classify generated text when models with identical architecture (including number of layers and hidden sizes) are trained on the same dataset. This allows us to assess the extent to which factors other than parameter size contribute to model-specific biases.
Specifically, we train seven 1B parameter LLMs with different learning rates, weight decay values, and random seeds. We confirmed that the trained models have comparable capabilities by verifying their downstream scores are nearly identical, as shown in Table 3 which shows the zero-shot performance of these LLMs on common sense reasoning tasks: PIQA (Bisk, Zellers et al. 2020), HellaSwag (Zellers, Holtzman et al. 2019), WinoGrande (Sakaguchi, Bras et al. 2021), ARC easy and challenge (Clark, Cowhey et al. 2018), and OpenBookQA (OBQA) (Mihaylov, Clark et al. 2018).
@@
Compared to the previous experiment, the Transformer-based classification accuracy decreases, suggesting that differences in model size (e.g., particularly in the capacity to capture the training data distribution) may contribute to the emergence of higher-order distinguishable patterns beyond simple word frequency differences.
@@ 4.4 What kind of Randomness during Training Induces Bias in LLMs?
To further investigate how these “natural fingerprints” develop throughout the training process, we examined whether model identifiability evolves predictably over time. We conduct additional experiments with models trained on 1T tokens instead of 100B tokens. We train three different Transformer-based classification models under both the Order setting (different training data order, same initialization) and the Init setting (different initialization, same training data orders), and perform classification experiments at six training checkpoints: 10B, 50B, 100B, 200B, 500B, 1T tokens (Figure 3).
@@
This convergent behavior suggests that despite differences in optimization paths caused by factors such as initialization or data order, models tend to exhibit their behavior when trained sufficiently on the same underlying dataset. However, in practical model development, it is often infeasible to perfectly control all training conditions except for the random seed and train models for sufficient steps. In other words, completely eliminating natural fingerprints is unlikely in real-world settings. Fortunately, the relatively small gap between accuracy of unigram-based and Transformer-based classifiers in Table 5 suggests that, when training data and model size are controlled, the distinguishability may largely be attributed to differences in word frequency distributions, which in turn makes the analysis of such differences more tractable.
4.5 Do Natural Fingerprints Emerge in Supervised Fine-Tuning?
Finally, we test whether natural fingerprints also emerge in smaller-scale training, such as supervised fine-tuning (SFT), where models start from pretrained weights, which may have a stronger influence on training dynamics compared to the random initialization used in pretraining.
We fine-tune the 1B pretrained models (trained on 100B tokens in the previous experiment) using the Alpaca dataset (Taori, Gulrajani et al. 2023). We use a learning rate of 2e-5 and train for one epoch with AdamW. Other hyperparameters are listed in Appendix A.